# A motivic Fundamental Lemma

### v3

@@ 4.2. Galois cohomology
We encode the fact that two cocycles $c,c’$ ’ are cohomologous if there exists a 1-coboundary $d$ such that $c=c’d$. Since this is a definable equivalence relation on 1-cocycles, this allows to consider $\mathrm{H}^{1}(\operatorname{Gal}(L/K),A)$ as an imaginary set (with parameters $(b,\tau)$).
@@ 4.4. Isotypical components
[…]
@@ 4.6. Quasi-split groups
Let $\rho$ be an $\operatorname{\mathrm{Out}}({\mathbb{G}})$-torsor over $X$ that is trivial over $\infty\in X$. By [22, Lemma 4.8] there is a finite étale cover $X^{\prime}$ of $X$ over which $\rho$ becomes trivial, and we can moreover assume that $X^{\prime}\to X$ is Galois. We then view $\rho$ as a morphism $\operatorname{\mathsf{Aut}}(X^{\prime}/X)\to\operatorname{\mathrm{Out}}({\mathbb{G}})$, which amounts to choose a finite number of points of $\operatorname{\mathrm{Out}}({\mathbb{G}})$ satisfying some conditions. By Section 4.2, if we fix an integer $r$ we get a definable set $\mathrm{H}^{1}(\operatorname{\mathsf{Aut}}(X^{\prime}/X),\operatorname{\mathrm{Out}}({\mathbb{G}}))$ representing the subset of $\operatorname{\mathrm{Out}}({\mathbb{G}})$-torsors on $X$ parametrized by morphisms $\operatorname{\mathsf{Aut}}(X^{\prime}/X)\to\operatorname{\mathrm{Out}}({\mathbb{G}})$ with $X^{\prime}$ a finite étale Galois cover of $X$ of degree $r$. Fix one such parameter $f$.
@@ 4.7. Higgs bundles
Given a split reductive group ${\mathbb{G}}$ on $k$, and a smooth projective curve $X$ with a $k$-point, we so far constructed a definable group ${\mathbf{G}}$. Fix in addition an even integer $d$ larger or equal to $22g-$ 2, where $g$ is the genus of $X$. To define Higgs bundles we need to fix a line bundle on $X$. Since the space of classes of line bundles of degree $d$ on $X$ is representable by a quasi-projective scheme $\mathrm{Pic}^{d}_{X/k}$, we add $\mathrm{Pic}^{d}_{X/k}$ to the set of parameters, and in what follows the definable set will depend on a parameter $D\in\mathrm{Pic}^{d}_{X/k}$ representing a line bundle on $X$ of degree $d$.
@@
To do so, first recall from Definition 3.3.1 that $\operatorname{\mathsf{A}}_{G_{K}}$ is the affine space of global sections $\mathrm{H}^{0}(X,\mathrm{c}_{D})$, and by [37, Lemme 4.13.1], its dimension (under the hypothesis $\deg(D)\geq 2g-2$) depends only on $g$, $\deg(D)$, and the rank and the number of roots of ${\mathbb{G}}$. Hence there is a definable set ${\mathbf{A}}$ such that ${\mathbf{A}}(K)=\operatorname{\mathsf{A}}_{G_{K}}(K)$ for every pseudo-finite field $K$. Using affine charts for $X$ provided by the projective embedding given by $D$, we can interpret elements of ${\mathbf{A}}$ as global sections hence for every pseudo-finite field $K$ and $a\in{\mathbf{A}}(K)$, $a$ determines a map $X(K)\to{\mathbf{c}}_{D}(K)$.
@@
(4.7.1) ${\widetilde{\operatorname{\mathsf{A}}}_{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{t}^{\mathrm{rs}}_{D,\infty}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathsf{A}}_{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{c}_{D,\infty}.}$
@@
Let $a\in\operatorname{\mathsf{A}}_{G_{K}}$. The cameral cover $\mathcal{C}_{a}$ is the $W$-equivariant morphism $\mathcal{C}_{a}\to X$, where $W$ is the twisted Weil group as in [37, 1.3.3], fitting into the cartesian diagram of $K$-schemes
[…]
@@
Recall from Section 3.5 that the anisotropic locus $\operatorname{\mathsf{A}}_{G_{K}}^{\mathrm{ani}}\subseteq\operatorname{\mathsf{A}}_{G_{K}}$ is the locus where the cameral curve is reduced and $\pi_{0}({\mathcal{P}}_{G_{K},a})$ is finite. By [37, Proposition 4.10.3], the second condition is equivalent to the fact that ${\mathfrak{t}_{D}}^{W_{\widetilde{a}}}=0$, where $\widetilde{a}=(a,\tilde{\infty})\in\tilde{\operatorname{\mathsf{A}}}_{G}$ is in the preimage of $a$ and $W_{\widetilde{a}}$ is the finite subgroup of $W\rtimes\operatorname{\mathrm{Out}}({\mathbb{G}})$ defined as follows, following [37, 1.3.6, 5.4.1].
@@
Since the degree $r$ is fixed, the cover $X^{\prime}$ is definable, and the open subset $U$ as well. Hence the above description of $W_{\widetilde{a}}$ depends definably on $\widetilde{a}\in\widetilde{{\mathbf{A}}}_{G}$. Hence the subset $\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}$ of $\widetilde{a}\in\widetilde{{\mathbf{A}}}_{G}$ such that $\mathcal{C}_{a}$ is reduced and ${\mathbf{t}}^{W_{\widetilde{a}}}=0$ is definable. The anisotropic locus ${{\mathbf{A}}}_{G}^{\mathrm{ani}}$ is the image of $\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}$ in ${\mathbf{A}}_{G}$, therefore definable as well. By [37, 6.1], $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$ is a non-empty open subset of $\operatorname{\mathsf{A}}_{G}$.
The Prym $\mathcal{P}_{G,a}$ acts on the fiber ${\mathcal{M}}_{G,a}$ and using the Kostant section [37, Section 4.2.4], is identified with an open subset of ${\mathcal{M}}_{G,a}$. Let $\mathcal{P}^{{\mathrm{ani}}}$ and ${\mathcal{M}}_{G}^{{\mathrm{ani}}}$ be the restriction of $\mathcal{P}_{G_{K}}$ and ${\mathcal{M}}_{G_{K}}$ to $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G_{K}}$. By [37, Proposition 6.1.3], $\mathcal{P}^{{\mathrm{ani}}}$ is a smooth separated Deligne-Mumford stack of finite type over $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G_{K}}$ and ${\mathcal{M}}_{G_{K}}^{{\mathrm{ani}}}$ is a smooth separated Deligne-Mumford stack of finite type over $K$. Moreover, the coarse moduli space $M_{G}^{{\mathrm{ani}}}$ of ${\mathcal{M}}_{G}^{{\mathrm{ani}}}$ is projective over $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$.
We claim that there exists a definable set ${\mathbf{M}}_{G}$ such that ${\mathbf{M}}_{G}^{\mathrm{ani}}(K)=M_{G}^{{\mathrm{ani}}}(K)$. By [37, Lemme 6.1.2], every $G$-Higgs bundle over $A^{{\mathrm{ani}}}$ is stable. Hence one can follow the construction by Faltings [16, Theorem II.5] of the coarse moduli space of stable $G$-Higgs bundles $M_{G}^{\mathrm{s}}$. This gives a quasi-projective embedding of $M_{G}^{\mathrm{s}}$ into some ${\mathbb{P}}^{N}\times X(G)^{*}$, where $N$ depends only on the dimension of $\operatorname{\mathsf{A}}_{G}$ and the given representation $\delta$ of ${\mathbb{G}}$, and $X(G)^{*}$ is the dual of the character lattice. The fibers over $X(G)^{*}$ are then definable, therefore the same is true for the restriction of $M_{G}^{\mathrm{s}}$ to the restriction to the anisotropic locus $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$. From the construction, one also gets that the stack structure is also definable in the sense of Definition 2.13.2.
@@
Let $\widetilde{{\mathbf{M}}}_{G}^{{\mathrm{ani}}}$ be the base change of ${\mathbf{M}}_{G}^{{\mathrm{ani}}}$ along $\widetilde{{\mathbf{A}}}_{G}\to{\mathbf{A}}_{G}$, which is then also definable. By Proposition [22, Prop. 4.38] there is a natural surjective morphism

### v4

@@ 4.2. Galois cohomology
We encode the fact that two cocycles $c,c$’ are cohomologous if there exists a 1-coboundary $d$ such that c=cdc=c’d. Since this is a definable equivalence relation on 1-cocycles, this allows us to consider $\mathrm{H}^{1}(\operatorname{Gal}(L/K),A)$ as an imaginary set (with parameters $(b,\tau)$).
@@ 4.4. Isotypical components
$0\longrightarrow\Gamma(K)\longrightarrow\Gamma(K^{\mathrm{alg}})\xrightarrow{\>1-\sigma\>}\Gamma(K^{\mathrm{alg}})\longrightarrow\mathrm{H}^{1}(K,\Gamma)\longrightarrow 0,$
@@ 4.6. Quasi-split groups
Let $\rho$ be an $\operatorname{\mathrm{Out}}({\mathbb{G}})$-torsor over $X$ that is trivial over $\infty\in X$. By [23, Lemma 4.8] there is a finite étale cover $X^{\prime}$ of $X$ over which $\rho$ becomes trivial, and we can moreover assume that $X^{\prime}\to X$ is Galois. We then view $\rho$ as a morphism $\operatorname{\mathsf{Aut}}(X^{\prime}/X)\to\operatorname{\mathrm{Out}}({\mathbb{G}})$, which amounts to choose a finite number of points of $\operatorname{\mathrm{Out}}({\mathbb{G}})$ satisfying some conditions. By Section 4.2, if we fix an integer $r$ we get a definable set $\mathrm{H}^{1}(\operatorname{\mathsf{Aut}}(X^{\prime}/X),\operatorname{\mathrm{Out}}({\mathbb{G}}))$ representing the subset of $\operatorname{\mathrm{Out}}({\mathbb{G}})$-torsors on $X$ parametrized by morphisms $\operatorname{\mathsf{Aut}}(X^{\prime}/X)\to\operatorname{\mathrm{Out}}({\mathbb{G}})$ with $X^{\prime}$ a finite étale Galois cover of $X$ of degree $r$. Fix one such parameter $f$.
@@ 4.7. Higgs bundles
Given a split reductive group $\mathbb{G}$} on $k$, and a smooth projective curve $X$ with a $k$-point, we so far constructed a definable group $\mathbf{G}$}. Fix in addition an even integer $d$ larger or equal to 2g22g-2, where $g$ is the genus of $X$. To define Higgs bundles we need to fix a line bundle on $X$. Since the space of classes of line bundles of degree $d$ on $X$ is representable by a quasi-projective scheme $\mathrm{Pic}^{d}_{X/k}$, we add $\mathrm{Pic}^{d}_{X/k}$ to the set of parameters, and in what follows the definable set will depend on a parameter $D\in\mathrm{Pic}^{d}_{X/k}$ representing a line bundle on $X$ of degree $d$.
@@
To do so, first recall from Definition 3.3.1 that $\operatorname{\mathsf{A}}_{G_{K}}$ is the affine space of global sections $\mathrm{H}^{0}(X,\mathrm{c}_{D})$, and by [38, Lemme 4.13.1], its dimension (under the hypothesis $\deg(D)\geq 2g-2$) depends only on $g$, $\deg(D)$, and the rank and the number of roots of $\mathbb{G}$}. Hence there is a definable set $\mathbf{A}$} such that ${\mathbf{A}}(K)=\operatorname{\mathsf{A}}_{G_{K}}(K)$ for every pseudo-finite field $K$. Using affine charts for $X$ provided by the projective embedding given by $D$, we can interpret elements of $\mathbf{A}$} as global sections hence for every pseudo-finite field $K$ and $\in{\mathbf{A}}(K), aa$ determines a map $X(K)\to{\mathbf{c}}_{D}(K)$.
@@
[…]
@@
Let $a\in\operatorname{\mathsf{A}}_{G_{K}}$. The cameral cover $\mathcal{C}_{a}$ is the $W$-equivariant morphism $\mathcal{C}_{a}\to X$, where $W$ is the twisted Weil group as in [38, 1.3.3], fitting into the cartesian diagram of $K$-schemes
(4.7.2) ${\mathcal{C}_{a}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{t}_{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a}$${\mathfrak{c}_{D}.}$
@@
Recall from Section 3.5 that the anisotropic locus $\operatorname{\mathsf{A}}_{G_{K}}^{\mathrm{ani}}\subseteq\operatorname{\mathsf{A}}_{G_{K}}$ is the locus where the cameral curve is reduced and $\pi_{0}({\mathcal{P}}_{G_{K},a})$ is finite. By [38, Proposition 4.10.3], the second condition is equivalent to the fact that ${\mathfrak{t}_{D}}^{W_{\widetilde{a}}}=0$, where $\widetilde{a}=(a,\tilde{\infty})\in\tilde{\operatorname{\mathsf{A}}}_{G}$ is in the preimage of $a$ and $W_{\widetilde{a}}$ is the finite subgroup of $W\rtimes\operatorname{\mathrm{Out}}({\mathbb{G}})$ defined as follows, following [38, 1.3.6, 5.4.1].
@@
Since the degree $r$ is fixed, the cover $X^{\prime}$ is definable, and the open subset $U$ as well. Hence the above description of $W_{\widetilde{a}}$ depends definably on $\widetilde{a}\in\widetilde{{\mathbf{A}}}_{G}$. Hence the subset $\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}$ of $\widetilde{a}\in\widetilde{{\mathbf{A}}}_{G}$ such that $\mathcal{C}_{a}$ is reduced and ${\mathbf{t}}^{W_{\widetilde{a}}}=0$ is definable. The anisotropic locus ${{\mathbf{A}}}_{G}^{\mathrm{ani}}$ is the image of $\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}$ in ${\mathbf{A}}_{G}$, therefore definable as well. By [38, 6.1], $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$ is a non-empty open subset of $\operatorname{\mathsf{A}}_{G}$.
The Prym $\mathcal{P}_{G,a}$ acts on the fiber ${\mathcal{M}}_{G,a}$ and using the Kostant section [38, Section 4.2.4], is identified with an open subset of ${\mathcal{M}}_{G,a}$. Let $\mathcal{P}^{{\mathrm{ani}}}$ and ${\mathcal{M}}_{G}^{{\mathrm{ani}}}$ be the restriction of $\mathcal{P}_{G_{K}}$ and ${\mathcal{M}}_{G_{K}}$ to $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G_{K}}$. By [38, Proposition 6.1.3], $\mathcal{P}^{{\mathrm{ani}}}$ is a smooth separated Deligne-Mumford stack of finite type over $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G_{K}}$ and ${\mathcal{M}}_{G_{K}}^{{\mathrm{ani}}}$ is a smooth separated Deligne-Mumford stack of finite type over $K$. Moreover, the coarse moduli space $M_{G}^{{\mathrm{ani}}}$ of ${\mathcal{M}}_{G}^{{\mathrm{ani}}}$ is projective over $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$.
We claim that there exists a definable set ${\mathbf{M}}_{G}$ such that ${\mathbf{M}}_{G}^{\mathrm{ani}}(K)=M_{G}^{{\mathrm{ani}}}(K)$. By [38, Lemme 6.1.2], every $G$-Higgs bundle over $A^{{\mathrm{ani}}}$ is stable. Hence one can follow the construction by Faltings [16, Theorem II.5] of the coarse moduli space of stable $G$-Higgs bundles $M_{G}^{\mathrm{s}}$. This gives a quasi-projective embedding of $M_{G}^{\mathrm{s}}$ into some ${\mathbb{P}}^{N}\times X(G)^{*}$, where $N$ depends only on the dimension of $\operatorname{\mathsf{A}}_{G}$ and the given representation $\delta$ of $\mathbb{G}$}, and $X(G)^{*}$ is the dual of the character lattice. The fibers over $X(G)^{*}$ are then definable, therefore the same is true for the restriction of $M_{G}^{\mathrm{s}}$ to the restriction to the anisotropic locus $\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G}$. From the construction, one also gets that the stack structure is also definable in the sense of Definition 2.13.2.
@@
Let $\widetilde{{\mathbf{M}}}_{G}^{{\mathrm{ani}}}$ be the base change of ${\mathbf{M}}_{G}^{{\mathrm{ani}}}$ along $\widetilde{{\mathbf{A}}}_{G}\to{\mathbf{A}}_{G}$, which is then also definable. By Proposition [23, Prop. 4.38] there is a natural surjective morphism
@@
4.8. Definability of moduli of $\beta$-twisted Higgs bundles
Recall from Section 3.6 that given a smooth projective curve $X$, $G$ a quasi-split reductive group on $X$ and $\beta$ a a $Z(G)$-gerbe on $X$, one can consider the moduli stack of $\beta$-twisted Higgs bundles ${\mathcal{M}}_{\beta}$, which is a smooth Deligne-Mumford stack over $A^{\mathrm{ani}}$.
Suppose now that we are in the situation of Section 4.7, with a smooth projective curve $X$ over $k$, a definable quasi-split reductive group $\mathbf{G}$} over $X$, and a line bundle $D$ considered as a point in the parameter space. Fix in addition a $Z({\mathbf{G}})$-gerbe $\beta$ represented by a 3-cocycle on a finite étale cover of $X$.
Proposition 4.8.1.
There is a definable Deligne-Mumford stack ${\mathbf{M}}_{G}^{\beta,{\mathrm{ani}}}$ such that for every pseudo-finite field $K$ extending $k$, ${\mathbf{M}}_{G}^{\beta,{\mathrm{ani}}}(K)=M_{G}^{\beta,{\mathrm{ani}}}(K)$.
Proof.
We first recall some constructions from [22].
Let $\bar{\mathbf{G}}={\mathbf{G}}_{\mathcal{E}}/Z({\mathbf{G}})$. From [22, Construction 2.11], consider the universal Higgs bundle $(E,\theta)$ on $X\times_{k}{\mathbb{M}}_{\bar{G}}$. Associate to it the $Z(G)$-gerbe $[E/G]$, this defines a global section $\beta^{\mathrm{univ}}\in\mathrm{H}^{0}({\mathbb{M}}_{\bar{G}},\mathrm{H}^{2}(X,\pi_{0}(Z(G))))$. Then let ${\mathbb{M}}_{\bar{G}}^{\beta}$ the largest open subset of ${\mathbb{M}}_{\bar{G}}$ where $\beta^{\mathrm{univ}}=\beta$. This yields a finite decomposition
(4.8.1) ${\mathbb{M}}_{\bar{G}}=\cup_{\beta\in\mathrm{H}^{2}(X,\pi_{0}(Z(G)))}{\mathbb{M}}_{\bar{G}}^{\beta}(K).$
By [22, Lemma 2.12], there is a canonical morphism
$F\colon{\mathbb{M}}_{G}^{\beta,{\mathrm{ani}}}\to{\mathbb{M}}_{\bar{G}}^{\beta,{\mathrm{ani}}}\times_{A_{\bar{G}}}A_{G},$
making ${\mathbb{M}}_{G}^{\beta,{\mathrm{ani}}}$ a $\mathrm{Sect}_{X}(BZ(G))$-torsor. Taking the rigidification of this morphism yields a morphism
$F\colon{\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}\to{\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}\times_{A_{\bar{G}}}A_{G},$
which is a $\pi_{0}(Z(G))$-torsor.
We now consider the definable case. From Proposition 4.7.1, there is a definable ${\mathbf{M}}_{\bar{G}}^{{\mathrm{ani}}}$. Taking the rigidification of (4.8.1), we get definable subsets ${\mathbf{M}}_{\bar{G}}^{\beta}\subset{\mathbf{M}}_{\bar{G}}$ since the condition $\beta^{\mathrm{univ}}=\beta$ is definable.
By [22, Lemma 2.12], ${\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}$ is a $\pi_{0}(Z(G))$-torsor over ${\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}\times_{A_{\bar{G}}}A_{G}$. Each $\pi_{0}(Z({\mathbf{G}}))$-torsor over ${\mathbf{M}}_{\bar{G}}^{\beta,{\mathrm{ani}}}\times_{{\mathbf{A}}_{\bar{G}}}{\mathbf{A}}_{G}$ is definable. To see that ${\mathbf{M}}_{G}^{\beta,{\mathrm{ani}}}$ is definable, is suffices then to see that the choice of the torsor is definable, i.e. that the assignment $K\mapsto F\in\mathrm{H}^{1}({\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}\times_{A_{\bar{G}}}A_{G},\pi_{0}(Z(G))$ is definable. But this amounts to check which twist of ${\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}$ has a section, which is a definable condition. ∎
Theorem 4.8.2.
For every definable gerbe $\beta$ on $X$, we have
$\chi_{\mathrm{Psf},\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{G}^{\beta,{\mathrm{ani}}},\mathrm{stab})=\chi_{\mathrm{Psf},\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{G}^{{\mathrm{ani}}},\mathrm{stab}).$
Proof.
By [22, Section 3.1], there exists a ${\mathcal{P}}_{G}$-torsor $\tau$ such that ${\mathcal{M}}_{G}^{\beta,\mathrm{reg}}$ is the twist of ${\mathcal{M}}_{G}^{\mathrm{reg}}$ by $\tau$, i.e. ${\mathcal{M}}_{G}^{\beta,\mathrm{reg}}\simeq{\mathcal{M}}_{G}^{\mathrm{reg},\tau}$, where $\operatorname{reg}$ means restriction to the regular locus.
As in the proof of the previous proposition, checking which twist is the correct one is a definable condition, hence given $\mathbf{G}$}, and a definable gerbe $\beta$, there is a definable ${\mathbf{P}}_{G}$-torsor $\tau$ and a definable bijection
${\mathbf{M}}_{G}^{\beta,\mathrm{reg}}\simeq{\mathbf{M}}_{G}^{\mathrm{reg},\tau}.$
This induces a definable bijection such that for every pseudo-finite field $K$,
$\widetilde{{\mathbf{M}}}_{G}^{\beta,{\mathrm{ani}},\natural}(K)\simeq\widetilde{{\mathbf{M}}}_{G}^{{\mathrm{ani}},\tau,\natural}(K),$
where as in Section 2.13, ${\mathbf{X}}^{\natural}(K)={\mathbf{X}}(K\mathopen{[\![}t\mathopen{]\!]})\cap{\mathbf{X}}^{\mathrm{reg}}(K\mathopen{(\!(}t\mathopen{)\!)})$.
We know that ${\mathbf{M}}_{G}^{\beta,{\mathrm{ani}}}$ is a definable Deligne-Mumford stack. So for every pseudo-finite field $K$, the map to the coarse moduli space
${\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}(K\mathopen{[\![}t\mathopen{]\!]})\to M_{G}^{\beta,{\mathrm{ani}}}(K\mathopen{[\![}t\mathopen{]\!]})$
is definable.
We now claim that $x\in M_{G}^{\beta,{\mathrm{ani}},\natural}(K)$ lifts to ${\mathcal{M}}_{G}^{\beta,{\mathrm{ani}}}(K\mathopen{[\![}t\mathopen{]\!]})$ if and only if its image in $M_{G}^{{\mathrm{ani}},\tau,\natural}(K)$ lifts to ${\mathcal{M}}_{G}^{{\mathrm{ani}},\tau}(K\mathopen{[\![}t\mathopen{]\!]})$. This is a first order property and when $K$ is a finite field is the statement of [22, Lemma 3.4]. So the property holds also for pseudo-finite fields. So the preimages of $\widetilde{{\mathbf{M}}}_{G}^{\beta}\subset I_{\hat{\mu}}\widetilde{{\mathbf{M}}}_{G}^{\beta}$ and $\widetilde{{\mathbf{M}}}_{G}^{\tau}\subset I_{\hat{\mu}}\widetilde{{\mathbf{M}}}_{G}^{\tau}$ under the specialization map $e$ (2.13.1) are identified.
Restricting to the Hitchin fiber over $a\in\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}$ we then obtain by the orbifold formula 2.13.4
$\widetilde{{\mathbf{M}}}^{\beta}_{G,a}]={\mathbb{L}}^{\dim\widetilde{{\mathbf{M}}}^{\beta}_{G,a}}\int_{e^{-1}(\widetilde{{\mathbf{M}}}^{\beta}_{G,a})}|\omega|={\mathbb{L}}^{\dim\widetilde{{\mathbf{M}}}^{\tau}_{G,a}}\int_{e^{-1}(\widetilde{{\mathbf{M}}}^{\tau}_{G,a})}|\omega|=[\widetilde{{\mathbf{M}}}^{\tau}_{G,a}].$
This equality holds in $\mathcal{C}_{\mathrm{Psf}}(\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}}\times S)$, where $S$ is the set of parameters. We then take its realization in $\mathrm{\bf K}(\operatorname{DA^{\acute{e}t}_{ct}}(\widetilde{{\mathbf{A}}}_{G}^{\mathrm{ani}},\mathrm{rel},\Lambda))$.
To obtain the comparison of the stable parts, we consider, for every $t$ in $H^{1}(K,\pi_{0}({\mathbf{P}}_{G,a}))$, the unramified twists $\left(\widetilde{{\mathbf{M}}}^{\beta}_{G,a}\right)^{t}$ and $\widetilde{{\mathbf{M}}}^{\tau+t}_{G,a}$ and repeat the above argument to obtain $\left[\left(\widetilde{{\mathbf{M}}}^{\beta}_{G,a}\right)^{t}\right]=[\widetilde{{\mathbf{M}}}^{\tau+t}_{G,a}]$. Taking realizations and summing up over all $t$ gives the result. ∎